# Representability of the direct sum of uniform $q$-matroids

### v1

@@ Introduction
Our first main result is the geometric characterization of the representation of the direct sum of uniform $q$-matroids. We prove that a $q$-system is a representation of such direct sum if and only if it satisfies some evasiveness properties; see Theorem 2.4. The crucial objects towards our characterization are the cyclic flats, which for uniform $q$-matroids are just the 00-subspace and the whole ground space.
@@ Acknowledgments
Gianira N. Alfarano is supported by the Swiss National Foundation through grant no. 210966 and by the FWO (Research Foundation Flanders) grant no. 1273624N. Alessandro Neri is supported by the FWO (Research Foundation Flanders) grant no. 12ZZB23N. Ferdinando Zullo is very grateful for the hospitality of the Ghent University, where he was a visiting researcher during the development of this research in September 2023. He is supported by the project COMBINE of “VALERE: VAnviteLli pEr la RicErca” of the University of Campania “Luigi Vanvitelli” and by Erasmus+ UE (Programme Countries) from Unicampania. This research is partially supported by the Italian National Group for Algebraic and Geometric Structures and their Applications (GNSAGA - INdAM).
@@ 1.1. Rank-metric codes and q-systems
An $[n,k]_{q^{m}/q}$ rank-metric code $\mathcal{C}$ is a $k$-dimensional $\mathbb{F}_{q^{m}}$-linear subspace of $\mathbb{F}_{q^{m}}^{n}$. The minimum rank distance of $\mathcal{C}$ is defined as
@@
If $d=\operatorname{d}_{\operatorname{r}}(\mathcal{C})$ is known, we write that $\mathcal{C}$ is an $[n,k,d]_{q^{m}/q}$ code. The parameters $n,m,k,d$ of an $[n,k,d]_{q^{m}/q}$ rank-metric code satisfy a Singleton-like bound [13], that reads as
@@
The support of $v$ is the column space of $\Gamma(v)$ for any basis $\Gamma$; we denote it by $\operatorname{supp}(v)$. The rank-weight of $v\in\mathbb{F}_{q^{m}}^{n}$ is the quantity $\operatorname{rk}(v)=\dim_{\mathbb{F}_{q}}(\operatorname{supp}(v))$. A generator matrix for an $[n,k,d]_{q^{m}/q}$ code $\mathcal{C}$ is a full-rank matrix $G\in\mathbb{F}_{q^{m}}^{k\times n}$ such that $\mathcal{C}=\operatorname{rowsp}(G)$. If the columns of one (and hence any) generator matrix of $\mathcal{C}$ are $\mathbb{F}_{q}$-linearly independent, then $\mathcal{C}$ is said to be nondegenerate. Two $[n,k,d]_{q^{m}/q}$ rank-metric codes $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are equivalent if and only if there exists $A\in\mathrm{GL}(n,\mathbb{F}_{q})$ such that $\mathcal{C}^{\prime}=\mathcal{C}A=\{vA:v\in\mathcal{C}\}$.
There is a geometric interpretation of $[n,k,d]_{q^{m}/q}$ rank-metric codes as $q$-systems.
@@
An $[n,k,d]_{q^{m}/q}$ system $\mathcal{S}$ is an $n$-dimensional $\mathbb{F}_{q}$-subspace of $\mathbb{F}_{q^{m}}^{k}$, such that $\langle\mathcal{S}\rangle_{\mathbb{F}_{q^{m}}}=\mathbb{F}_{q^{m}}^{k}$ and
@@
Moreover, two $[n,k,d]_{q^{m}/q}$ systems $\mathcal{S}$ and $\mathcal{S}^{\prime}$ are equivalent if there exists an $\mathbb{F}_{q^{m}}$-isomorphism $\varphi\in\operatorname{GL}(k,\mathbb{F}_{q^{m}})$ such that
@@
Let $\mathcal{S}$ be an $[n,k]_{q^{m}/q}$ system. For each $\mathbb{F}_{q}$-subspace $V$ of $\mathbb{F}_{q^{m}}^{k}$, we define the weight of $V$ in $\mathcal{S}$ as the integer
@@
In [32, 29] it has been proved that there is a one-to-one correspondence between equivalence classes of nondegenerate $[n,k,d]_{q^{m}/q}$ codes and equivalence classes of $[n,k,d]_{q^{m}/q}$ systems and the correspondence can be explained as follows. Let $\mathcal{C}$ be an $[n,k,d]_{q^{m}/q}$ code and $G$ be a generator matrix for $\mathcal{C}$. Define $\mathcal{S}$ to be the $\mathbb{F}_{q}$-span of the columns of $G$. In this case $\mathcal{S}$ is also called a $q$-system associated with $\mathcal{C}$. Vice versa, given an $[n,k,d]_{q^{m}/q}$ system $\mathcal{S}$, define $G$ to be the matrix whose columns are an $\mathbb{F}_{q}$-basis of $\mathcal{S}$ and let $\mathcal{C}$ be the code generated by $G$. $\mathcal{C}$ is also called a code associated with $\mathcal{S}$. For more details on this correspondence, see also [2].
Finally recall the following result which provides a natural description of the supports of codewords of an $[n,k]_{q^{m}/q}$ code $\mathcal{C}$ in terms of a $q$-system $\mathcal{S}$ associated to $\mathcal{C}$.
@@
Let $\mathcal{C}$ be an $[n,k]_{q^{m}/q}$ non-degenerate rank-metric code and let $\mathcal{S}$ be the $\mathbb{F}_{q}$-span of the columns of a generator matrix $G$. Consider the isomorphism
@@
Let $\mathcal{A}$ be a collection of $\mathbb{F}_{q}$-subspaces of $\mathbb{F}_{q^{m}}^{k}$, let $\mathcal{S}$ be an $[n,k]_{q^{m}/q}$ system and let $h$ be a positive integer. We say that $\mathcal{S}$ is $(\mathcal{A},h)$-evasive if
@@
Let $\mathcal{S}$ be an $[n,k]_{q^{m}/q}$ system, and let $h,r$ be positive integers such that $1\leq h\leq k$. We say that $\mathcal{S}$ is $(h,r)$-evasive if $\mathcal{S}$ is $(\Lambda_{h},r)$-evasive. When $h=r$, we say that $\mathcal{S}$ is $h$-scattered. Finally, a $1$-scattered $q$-system will be simply called scattered.
The study of scattered $q$-systems originated in [6]. This notion was then generalized to any $h$ in [12], although the $(k-1)$-scatteredness was previously used in [33] due to its connection with MRD codes. The complete correspondence between maximum $h$-scattered systems and MRD codes was then highlighted in [35]. More recently, the more general property of evasiveness was analyzed in [4] and [21]. The connection between evasiveness of $q$-systems and generalized rank weights of a corresponding code can be instead found in [24].
@@
that is the $\mathbb{F}_{q^{m}}$-dimension of the $\mathbb{F}_{q^{m}}$-span of any $\mathbb{F}_{q}$-basis of $V$. When we consider $\rho$ restricted to the $\mathbb{F}_{q}$-subspaces of an $[n,k]_{q^{m}/q}$ system $\mathcal{S}$, we will write $\rho_{\mathcal{S}}$.
@@
Given a vector space $E$, we denote by $\mathcal{L}(E)$ the lattice of all subspaces of $E$. For any direct sum $E=E_{1}\oplus E_{2}\oplus\cdots\oplus E_{t}$ of $\mathbb{F}_{q}$-vector spaces $E_{1},E_{2},\ldots,E_{t}$, we denote by $\pi_{i}:E\longrightarrow E_{i}$ the corresponding projections and by $\iota_{i}:E_{i}\longrightarrow E$ the canonical embedding. Furthermore, for $n_{1},n_{2},\ldots,n_{t}\in\mathbb{N}$ and $n=n_{1}+n_{2}+\cdots+n_{t}$, we set $\mathbb{F}_{q}^{n}=\mathbb{F}_{q}^{n_{1}}\oplus\mathbb{F}_{q}^{n_{2}}\oplus\cdots\oplus\mathbb{F}_{q}^{n_{t}}$ such that $\pi_{i}:\mathbb{F}_{q}^{n}\longrightarrow\mathbb{F}_{q}^{n_{i}}$ is the projections onto the corresponding $n_{i}$ coordinates. Let $Q$ be a prime power and define $\iota_{i}:\mathcal{L}(\mathbb{F}_{Q}^{n_{i}})\longrightarrow\mathcal{L}(\mathbb{F}_{Q}^{n})$ such that for any $X\in\mathcal{L}(\mathbb{F}_{Q}^{n_{i}})$ then
@@
Let $\mathcal{S}_{1},\ldots,\mathcal{S}_{t}$ be $[n_{i},k_{i}]_{q^{m}/q}$ systems, for $1\leq i\leq t$. Then, denote by $\bigoplus_{i=1}^{t}\mathcal{S}_{i}$ the $\left[\sum_{i=1}^{t}n_{i},\sum_{i=1}^{t}k_{i}\right]_{q^{m}/q}$ system given by $\iota_{1}(\mathcal{S}_{1})+\ldots+\iota_{t}(\mathcal{S}_{t})$.
@@
There is also an equivalent geometric interpretation of a representable $q$-matroid in terms of $q$-systems. This was proved in [3, Theorem 4.6], whose original statement may look slightly different. However, it is equivalent to the following reformulation.
@@
Let $G$ be the generator matrix of a nondegenerate $[n,k]_{q^{m}/q}$ code, and let $\mathcal{S}$ be any $[n,k]_{q^{m}/q}$ system associated to it. Them $\mathcal{M}_{G}\cong(\mathcal{S},\rho_{\mathcal{S}})$.
In other words, Theorem 1.11 characterizes representable $q$-matroids as those coming from a $q$-system. More precisely, a $q$-matroid $\mathcal{M}$ of rank $k$ and height $n$ without loops is $\mathbb{F}_{q^{m}}$-representable if and only if it is equivalent to a $q$-matroid $(\mathcal{S},\rho_{\mathcal{S}})$, for some $[n,k]_{q^{m}/q}$ system $\mathcal{S}$.
@@
Let $\mathcal{U}_{k,n}(q)$ be the uniform $q$-matroid of rank $k$ with ground space $\mathbb{F}_{q}^{n}$. Then the trivial uniform $q$-matroid $\mathcal{U}_{0,n}(q)$ and the free uniform $q$-matroid $\mathcal{U}_{n,n}$ are representable over $\mathbb{F}_{q}$. In particular, $\mathcal{U}_{0,n}(q)$ is represented by the $1\times n$ zero matrix and $\mathcal{U}_{n,n}(q)$ by the $n\times n$ identity matrix. For $0<k<n$, the uniform $q$-matroid $\mathcal{U}_{k,n}(q)$ is representable over $\mathbb{F}_{q^{m}}$ if and only if $m\geq n$. Indeed, a matrix $G\in\mathbb{F}_{q^{m}}^{k\times n}$ represents $\mathcal{U}_{k,n}(q)$ if and only if $\operatorname{rk}(GY^{\top})=k$ for all $Y\in\mathbb{F}_{q}^{k\times n}$ of rank $k$. But this is equivalent to $G$ generating an MRD code and such a matrix $G$ exists if and only if $m\geq n$; see e.g. [13]. Analogously, in the language of $q$-systems, a $\mathbb{F}_{q^{m}}$-representation of $\mathcal{U}_{k,n}(q)$ is given by any $[n,k]_{q^{m}/q}$ system which is $(k-1)$-scattered, and this is known to exist if and only if $m\geq n$; see e.g. [33].
Like in the classical case, not all $q$-matroids are representable; see for instance the Vámos $q$-matroid defined in [8, Example 17]. However, so far, very little is known about representability of $q$-matroids.

### v2

@@ Introduction
Our first main result is the geometric characterization of the representation of the direct sum of uniform $q$-matroids. We prove that a $q$-system is a representation of such direct sum if and only if it satisfies some evasiveness properties; see Theorem 2.4. The crucial objects towards our characterization are the cyclic flats, which for uniform $q$-matroids are just the 0-subspace and the whole ground space.
@@ Acknowledgments
Gianira N. Alfarano was supported by the Swiss National Foundation through grant no. 210966 and by the FWO (Research Foundation Flanders) grant no. 1273624N. Alessandro Neri was supported by the FWO (Research Foundation Flanders) grant no. 12ZZB23N and by the INdAM - GNSAGA Project CUP E53C24001950001 “Noncommutative polynomials in coding theory”. Ferdinando Zullo is very grateful for the hospitality of Ghent University, where he was a visiting researcher during the development of this research in September 2023. He is supported by the project COMBINE of the University of Campania “Luigi Vanvitelli” and by Erasmus+ UE (Programme Countries) from Unicampania. This research is partially supported by the Italian National Group for Algebraic and Geometric Structures and their Applications (GNSAGA - INdAM).
@@ 1.1. Rank-metric codes and q-systems
An $n,k]_{q^{m}/q}$ rank-metric code $\mathcal{C}$ is a $k$-dimensional $\mathbb{F}_{q^{m}}$-linear subspace of $\mathbb{F}_{q^{m}}^{n}$. The minimum rank distance of $\mathcal{C}$ is defined as
@@
If $d=\operatorname{d}_{\operatorname{r}}(\mathcal{C})$ is known, we write that $\mathcal{C}$ is an $n,k,d]_{q^{m}/q}$ code. The parameters $n,m,k,d$ of an $n,k,d]_{q^{m}/q}$ rank-metric code satisfy a Singleton-like bound [13], that reads as
@@
The support of $v$ is the column space of $\Gamma(v)$ for any basis $\Gamma$; we denote it by $\operatorname{supp}(v)$. The rank-weight of $v\in\mathbb{F}_{q^{m}}^{n}$ is the quantity $\operatorname{rk}(v)=\dim_{\mathbb{F}_{q}}(\operatorname{supp}(v))$. A generator matrix for an $n,k,d]_{q^{m}/q}$ code $\mathcal{C}$ is a full-rank matrix $G\in\mathbb{F}_{q^{m}}^{k\times n}$ such that $\mathcal{C}=\operatorname{rowsp}(G)$. If the columns of one (and hence any) generator matrix of $\mathcal{C}$ are $\mathbb{F}_{q}$-linearly independent, then $\mathcal{C}$ is said to be nondegenerate. Two $n,k,d]_{q^{m}/q}$ rank-metric codes $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are equivalent if and only if there exists $A\in\mathrm{GL}(n,\mathbb{F}_{q})$ such that $\mathcal{C}^{\prime}=\mathcal{C}A=\{vA:v\in\mathcal{C}\}$.
There is a geometric interpretation of $n,k,d]_{q^{m}/q}$ rank-metric codes as $q$-systems.
@@
An $n,k,d]_{q^{m}/q}$ (or $n,k]_{q^{m}/q}$) system $\mathcal{S}$ is an $n$-dimensional $\mathbb{F}_{q}$-subspace of $\mathbb{F}_{q^{m}}^{k}$, such that $\langle\mathcal{S}\rangle_{\mathbb{F}_{q^{m}}}=\mathbb{F}_{q^{m}}^{k}$ and
@@
The full $q$-system of $\mathbb{F}_{q^{m}}^{k}$ is the $mk,k]_{q^{m}/q}$ system $\mathcal{S}=\mathbb{F}_{q^{m}}^{k}$. Moreover, two $n,k,d]_{q^{m}/q}$ systems $\mathcal{S}$ and $\mathcal{S}^{\prime}$ are equivalent if there exists an $\mathbb{F}_{q^{m}}$-isomorphism $\varphi\in\operatorname{GL}(k,\mathbb{F}_{q^{m}})$ such that
@@
Let $\mathcal{S}$ be an $n,k]_{q^{m}/q}$ system. For each $\mathbb{F}_{q}$-subspace $V$ of $\mathbb{F}_{q^{m}}^{k}$, we define the weight of $V$ in $\mathcal{S}$ as the integer
@@
In [32, 29] it has been proved that there is a one-to-one correspondence between equivalence classes of nondegenerate $n,k,d]_{q^{m}/q}$ codes and equivalence classes of $n,k,d]_{q^{m}/q}$ systems and the correspondence can be explained as follows. Let $\mathcal{C}$ be an $n,k,d]_{q^{m}/q}$ code and $G$ be a generator matrix for $\mathcal{C}$. Define $\mathcal{S}$ to be the $\mathbb{F}_{q}$-span of the columns of $G$; it turns out to be an $n,k,d]_{q^{m}/q}$ system. In this case $\mathcal{S}$ is also called a $q$-system associated with $\mathcal{C}$. Vice versa, given an $n,k,d]_{q^{m}/q}$ system $\mathcal{S}$, define $G$ to be the matrix whose columns are an $\mathbb{F}_{q}$-basis of $\mathcal{S}$ and let $\mathcal{C}$ be the code generated by $G$; it turns out to be an $n,k,d]_{q^{m}/q}-$code. $\mathcal{C}$ is also called a code associated with $\mathcal{S}$. For more details on this correspondence, see also [2].
Finally recall the following result which provides a natural description of the supports of codewords of an $n,k]_{q^{m}/q}$ code $\mathcal{C}$ in terms of a $q$-system $\mathcal{S}$ associated to $\mathcal{C}$.
@@
Let $\mathcal{C}$ be an $n,k]_{q^{m}/q}$ non-degenerate rank-metric code and let $\mathcal{S}$ be the $\mathbb{F}_{q}$-span of the columns of a generator matrix $G$. Consider the isomorphism
@@
Let $\mathcal{A}$ be a collection of $\mathbb{F}_{q}$-subspaces of $\mathbb{F}_{q^{m}}^{k}$, let $\mathcal{S}$ be an $n,k]_{q^{m}/q}$ system and let $h$ be a positive integer. We say that $\mathcal{S}$ is $(\mathcal{A},h)$-evasive if
@@
Let $\mathcal{S}$ be an $n,k]_{q^{m}/q}$ system, and let $h,r$ be positive integers such that $1\leq h\leq k$. We say that $\mathcal{S}$ is $(h,r)$-evasive if $\mathcal{S}$ is $(\Lambda_{h},r)$-evasive. When $h=r$, we say that $\mathcal{S}$ is $h$-scattered. Finally, a $1$-scattered $q$-system will be simply called scattered.
The study of scattered $q$-systems originated in [6]. This notion was then generalized to any $h$ in [12], although the (k1)(k-1)-scatteredness was previously used in [33] due to its connection with MRD codes. The complete correspondence between maximum $h$-scattered systems and MRD codes was then highlighted in [35]. More recently, the more general property of evasiveness was analyzed in [4] and [21]. The connection between evasiveness of $q$-systems and generalized rank weights of a corresponding code can be instead found in [24].
@@
that is the $\mathbb{F}_{q^{m}}$-dimension of the $\mathbb{F}_{q^{m}}$-span of any $\mathbb{F}_{q}$-basis of $V$. When we consider $\rho$ restricted to the $\mathbb{F}_{q}$-subspaces of an $n,k]_{q^{m}/q}$ system $\mathcal{S}$, we will write $\rho_{\mathcal{S}}$.
@@
Given a vector space $E$, we denote by $\mathcal{L}(E)$ the lattice of all subspaces of $E$. For any direct sum $E=E_{1}\oplus E_{2}\oplus\cdots\oplus E_{t}$ of $\mathbb{F}_{q}$-vector spaces $E_{1},E_{2},\ldots,E_{t}$, we denote by $\pi_{i}:E\longrightarrow E_{i}$ the corresponding projections and by $\iota_{i}:E_{i}\longrightarrow E$ the canonical embedding. Furthermore, for $n_{1},n_{2},\ldots,n_{t}\in\mathbb{N}$ and $n=n_{1}+n_{2}+\cdots+n_{t}$, we set $\mathbb{F}_{q}^{n}=\mathbb{F}_{q}^{n_{1}}\oplus\mathbb{F}_{q}^{n_{2}}\oplus\cdots\oplus\mathbb{F}_{q}^{n_{t}}$ such that $\pi_{i}:\mathbb{F}_{q}^{n}\longrightarrow\mathbb{F}_{q}^{n_{i}}$ is the projection onto the corresponding $n_{i}$ coordinates. Let $Q$ be a prime power and define $\iota_{i}:\mathcal{L}(\mathbb{F}_{Q}^{n_{i}})\longrightarrow\mathcal{L}(\mathbb{F}_{Q}^{n})$ such that for any $X\in\mathcal{L}(\mathbb{F}_{Q}^{n_{i}})$ then
@@
Let $\mathcal{S}_{1},\ldots,\mathcal{S}_{t}$ be $n_{i},k_{i}]_{q^{m}/q}$ systems, for $1\leq i\leq t$. Then, denote by $\bigoplus_{i=1}^{t}\mathcal{S}_{i}$ the $\left[\sum_{i=1}^{t}n_{i},\sum_{i=1}^{t}k_{i}\right]_{q^{m}/q}$ system given by $\iota_{1}(\mathcal{S}_{1})+\ldots+\iota_{t}(\mathcal{S}_{t})$.
@@
There is also an equivalent geometric interpretation of a representable $q$-matroid in terms of $q$-systems. This was proved in [3, Theorem 4.6], whose original statement may look slightly different. Indeed, there it was proved by using the independent spaces. However, it is equivalent to the following reformulation, which uses the notion of rank function.
@@
Let $G$ be the generator matrix of a nondegenerate $n,k]_{q^{m}/q}$ code, and let $\mathcal{S}$ be any $n,k]_{q^{m}/q}$ system associated to it. Then $\mathcal{M}_{G}\cong(\mathcal{S},\rho_{\mathcal{S}})$.
Proof.
Let us consider the isomorphism of $\mathbb{F}_{q}$-vector spaces
$\psi_{G}:\mathbb{F}_{q}^{n}\longrightarrow\mathcal{S},\quad v\longmapsto vG^{\top}.$
For any given subspace $U\in\mathcal{L}(\mathbb{F}_{q}^{n})$, let $A^{U}$ be a matrix whose columns form a basis of $U$. Then,
$\rho_{\mathcal{S}}(\psi_{G}(U))=\dim_{\mathbb{F}_{q^{m}}}\langle\psi_{G}(U)\rangle_{\mathbb{F}_{q^{m}}}=\dim_{\mathbb{F}_{q^{m}}}(\mathrm{colsp}(GA^{U}))=\operatorname{rk}(GA^{U})=\rho(U),$
where $\rho$ is the rank function defined in (3) of the $q$-matroid $\mathcal{M}_{G}$. ∎
In other words, Theorem 1.11 characterizes representable $q$-matroids as those coming from a $q$-system. More precisely, a $q$-matroid $\mathcal{M}$ of rank $k$ and height $n$ without loops is $\mathbb{F}_{q^{m}}$-representable if and only if it is equivalent to a $q$-matroid $(\mathcal{S},\rho_{\mathcal{S}})$, for some $n,k]_{q^{m}/q}$ system $\mathcal{S}$. Therefore, $(\mathcal{S},\rho_{\mathcal{S}})$ is called an $\mathbb{F}_{q^{m}}$-representation of $\mathcal{M}_{G}$.
@@
Let $\mathcal{U}_{k,n}(q)$ be the uniform $q$-matroid of rank $k$ with ground space $\mathbb{F}_{q}^{n}$. Then the trivial uniform $q$-matroid $\mathcal{U}_{0,n}(q)$ and the free uniform $q$-matroid $\mathcal{U}_{n,n}$ are representable over $\mathbb{F}_{q}$. In particular, $\mathcal{U}_{0,n}(q)$ is represented by the $1\times n$ zero matrix and $\mathcal{U}_{n,n}(q)$ by the $n\times n$ identity matrix. For $0<k<n$, the uniform $q$-matroid $\mathcal{U}_{k,n}(q)$ is representable over $\mathbb{F}_{q^{m}}$ if and only if $m\geq n$. Indeed, a matrix $G\in\mathbb{F}_{q^{m}}^{k\times n}$ represents $\mathcal{U}_{k,n}(q)$ if and only if $\operatorname{rk}(GY^{\top})=k$ for all $Y\in\mathbb{F}_{q}^{k\times n}$ of rank $k$. But this is equivalent to $G$ generating an MRD code and such a matrix $G$ exists if and only if $m\geq n$; see e.g. [19, Ex. 2.4]. Analogously, in the language of $q$-systems, a $\mathbb{F}_{q^{m}}$-representation of $\mathcal{U}_{k,n}(q)$ is given by any $n,k]_{q^{m}/q}$ system which is (k1)(k-1)-scattered, and this is known to exist if and only if $m\geq n$; see e.g. [33].
Like in the classical case, not all $q$-matroids are representable; see for instance the Vámos $q$-matroid defined in [8, Example 17] , as shown in [16]. However, so far, very little is known about representability of $q$-matroids.